# PhishLumos: An Adaptive Multi-Agent System for Proactive Phishing Campaign Mitigation

### v1

@@ IV-B Proactive Mitigation and Social Impact
PhishLumos demonstrated excellent performance, achieving a median campaign coverage of 100%. This indicates that the generated rules were effective at identifying nearly all known URLs within a given campaign. The system showed powerful proactive detection capabilities. It discovered 77,391 URLs not present in our original ground truth data. We evaluated these new URLs using VirusTotal [22], which uses results from 90 security engines. Of these URLs, 62,321 (80.5%) were later flagged as malicious by at least one engine. This high confirmation rate is conservative, as our method finds cloaked sites before they are blocklisted. The fact that nearly 20% of these newly discovered URLs still evaded this extensive panel of scanners underscores the sophistication of the threats PhishLumos uncovers. To ensure practical safety, a parallel check against top-ranking websites [24] confirmed zero false positives, underscoring the precision of our method. On average, this translates to discovering 751 new malicious URLs per campaign.

### v2

@@ IV-B Proactive Mitigation and Social Impact
PhishLumos demonstrated excellent performance, achieving a median campaign coverage of 100%. This indicates that the generated rules were effective at identifying nearly all known URLs within a given campaign. The system showed powerful proactive detection capabilities. It discovered 77,391 URLs not present in our original ground truth data. We evaluated these new URLs using VirusTotal [19], which uses results from 90 security engines. Of these URLs, 62,321 (80.5%) were later flagged as malicious by at least one engine. This high confirmation rate is conservative, as our method finds cloaked sites before they are blocklisted. The fact that nearly 20% of these newly discovered URLs still evaded this extensive panel of scanners underscores the sophistication of the threats PhishLumos uncovers. To ensure practical safety, a parallel check against top-ranking websites [22] confirmed zero false positives, underscoring the precision of our method. On average, this translates to discovering 751.4 new malicious URLs per campaign.